OpenStudy (anonymous):

Solve the equation x − 3 = 13 for x. Please show your work and solution in the space provided.

jimthompson5910 (jim_thompson5910):

We're subtracting 3 from x, so we want to undo this by adding 3 to both sides. This will isolate x x − 3 = 13 x − 3+3 = 13+3 x + 0 = 16 x = 16 So the solution is x = 16

OpenStudy (anonymous):

Thank You Love!(:

jimthompson5910 (jim_thompson5910):

you're welcome, let me know if you have any questions about it